# SCANIA Component X Dataset: A Real-World Multivariate Time Series Dataset for Predictive Maintenance

### v1

@@ 3.3 Test set
Specification information of 5045 test vehicles is collected in test_specifications.csv file containing categorical values varying between Cat0, Cat1,…, and Cat8. Furthermore, this file has no missing values.
@@ 4 Technical Validation
OEMs usually hesitate to share their device operational data for many reasons, such as GDPR compliance, confidentiality agreements (NDAs), ownership of data, device failure rates, etc. As a result, many works on data-driven PdM are limited to private datasets [18, 19, 20, 21, 22, 23]. This prevents researchers who worked with real data comparing their methods performances with each other. The newly shared dataset by SCANIA can serve as a benchmark in the field, making it easier for future researchers to compare their methods and generate reproducible results.
The proposed dataset is highly conducive to various machine learning tasks like regression, anomaly detection, survival analysis, and classification, in the PdM field. This merit arises from two supreme characteristics: being a real-world dataset collected from actual trucks and exhibiting a multi-variate time series structure. In regression tasks, this dataset is valuable for estimating components’ remaining useful life or predicting the time until the next repair. Leveraging historical and time-to-event data, regression models can provide insightful estimates. Survival analysis is a powerful technique that predicts the survival function and the probability of an event happening at a specific time while considering censored data in the training phase. The blue vehicle populations in Fig 4 are referred to as censored data. The temporal structure of the dataset is crucial for the effective application of survival analysis. Furthermore, for classification tasks, using time series data, the model can classify whether a vehicle is going to fail within a specific time window or not.
To evaluate the performance of the aforementioned models, a cost function is suggested by experts of the company, which is defined by the sum of the different “Cost_n_m” multiplied by the number of instances, resulting in a summarized cost (Total_cost).
@@
5 Data availability
The proposed dataset collected from component X of SCANIA trucks is publicly available and licensed under CC BY 4.0[24]. This dataset consists of seven individual data files, all in CSV (Comma-Separated Values) format. Please note that the data file containing the labels for the test set is currently unavailable since this dataset was released for the ongoing IDA 2024 industrial challenge. The goal is to estimate these labels for the test set, and the corresponding data file will be published when the industrial challenge is concluded.

### v2

@@ 3.3 Testing set
3.3.2 test_labels.csv
This data file includes the class label for 5045 vehicles in the test set that their last readouts are randomly selected in five classes of 0, 1, 2, 3, and 4. Like the validation_labels, this data file is also imbalanced and is skewed toward class 0, which contains 4903 samples. In comparison, classes 1, 2, 3, and 4 include 26, 15, 41, and 60 samples.
@@
Specification information of 5045 test vehicles is collected in the test_specifications.csv file containing eight categorical features with values varying between Cat0, Cat1,…, and Cat28, with no missing values. Figure 6 illustrates the categories of each feature in the (*_specifications.csv files, including for the test set. Compared to the bars related to the train_specifications and validation_specifications, the red bars representing the test_specification exhibit the same normalized distribution, indicating the consistency in category distribution across all three data parts.
@@ 4 Technical Validation
OEMs usually hesitate to share their device operational data for many reasons, such as GDPR compliance, confidentiality agreements (NDAs), ownership of data, device failure rates, etc. As a result, many works on data-driven PdM are limited to private datasets 26, 27, 28, 29, 30, 31. This prevents researchers who worked with real data comparing their methods performances with each other. The newly shared dataset 5 by SCANIA can serve as a benchmark in the field, making it easier for future researchers to compare their methods and generate reproducible results.
The proposed dataset 5 is highly conducive to various machine learning tasks like regression, anomaly detection, survival analysis, and classification, in the PdM field. This merit arises from two supreme characteristics: being a real-world dataset collected from actual trucks and exhibiting a multi-variate time series structure. In regression tasks, this dataset 5 is valuable for estimating components’ remaining useful life or predicting the time until the next repair. Leveraging historical and time-to-event data, regression models can provide insightful estimates. Survival analysis is a powerful technique that predicts the survival function and the probability of an event happening at a specific time while considering censored data in the training phase. The blue vehicle populations in Fig. 4 are referred to as censored data. The temporal structure of the dataset is crucial for the effective application of survival analysis. Furthermore, for classification tasks, using time series data, the model can classify whether a vehicle is going to fail within a specific time window or not.
To evaluate the performance of the aforementioned models, a cost function is suggested by experts of the company, which is defined by the sum of the different “Cost_$\_n\_m$” multiplied by the number of instances, resulting in a summarized cost (Total_cost).